OpenStudy (anonymous):

which function represents the linear pattern shown it the table. x f(x) 1 -2 2 1 3 4 4 7

OpenStudy (anonymous):

the options are a. f(x)= 3x-5 b. f(x)=2x-4 c. f(x)=x+3 d.f(x)=x+3

OpenStudy (anonymous):

you can find the slope simply by counting how many units y increases with respect to x i.e. if x increases by one how many units does y increase? Slope or (m)= rise/run and the equation of the line is y=-mx+b

OpenStudy (anonymous):

in my mind I approach this, as I would go about an arithmetic sequence. Since it is linear, it would be an arithmetic sequence not geometric. so, lets write it out -2, 1, 4, 7 -2 + __ = 1 1 + __ = 4 4 + __ = 7 it is obviously increasing, but by what value? That value is the SLOPE.

OpenStudy (anonymous):

ok now wat

OpenStudy (anonymous):

Tell me by how much is each "term" in the f(x) increasing ?

OpenStudy (anonymous):

its -2,1,4,and then 7

OpenStudy (anonymous):

yes, and by how much is each term increasing ?

OpenStudy (anonymous):

about 2

OpenStudy (anonymous):

-2 + x = 1 1 + x = 4 4 + x= 7 solve for x in each of the equations I am posting above.

OpenStudy (anonymous):

i dont get the first problem but 1+3=4 4+3=7

OpenStudy (anonymous):

yEs, so it goes up by 3, that means the slope is 3.